246 Count the valence electrons in the molecule CH20. Which atom is the central atom? How did you decide? Draw its Lewis structure. 5. 6. Count the valence electrons in the molecule H2Se. Which atom is the central atom? How did you decide? Draw its Lewis structure.
Answer #1

Solution :-

Q5) CH2O

Total valence electrons = (1*C)+(2*H)+(1*O)

                                            = (1*4)+(2*1)+(1*6)

                                            = 12

The valency of the carbon is 4 while the valency of the oxygen is 2 , hydrogen can only form 1 bond each

Therefore the central atom is carbon because it can form more bonds among the elements in the formula

Lewis structure is drawn in the following image

Q6) H2Se

Total valence electrons = (2*H)+(1*Se)

                                          = (2*1)+(1*6)

                                           = 2

Central atom is Se because it is group 6 element which can form two bonds while the H only makes single bond.

Following image shows the lewis structure.

Se--r rt V>